Two charged particles are enclosed in an infinite cylindrical container. The
charge of the first particle is +20 nC and it is located 35 cm away from the other
charged particle. Calculate the charge of the second particle if the net electric
force of the system is equal to 2.83x10-5 N and is attractive.
1
Expert's answer
2020-12-14T07:20:13-0500
"F=k\frac{qQ}{r^2}\\2.83\cdot10^{-5}=-(8.99\cdot10^{9})\frac{(20\cdot10^{-9})Q}{0.35^2}\\
Q=-19.3\cdot10^{-9}C=-19.3\ nC"